OpenStudy (secret-ninja):

Did I do this right? Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. y - 6 = -3x - 3 y - 6 = -3(x + 1) y = -3x + 3 3x + y = 3

OpenStudy (secret-ninja):

I think its B...

OpenStudy (texaschic101):

y - y1 = m(x - x1) slope(m) = -3 (-1,6)...x1 = -1 and y1 = 6 now we sub y - 6 = -3(x -(-1) y - 6 = -3(x + 1) yep..its B

OpenStudy (secret-ninja):

Thank you!

OpenStudy (texaschic101):

no problem :)